HOW-TOs Guides huge Getting Started Guide Red Hat Reference Guide Red Hat Customisation Guide mailing list subscribe How To Ask Questions The Smart Way

pinfo
info

$

$ info chmod

n p u l h

emacs
info
rpm
httpd
man
httpd
rpm

q

httpd

command | | httpd l d c i v

httpd
/etc/passwd

f

man
rpm

Red Hat Linux Customization Guide Maximum RPM

dpkg
apt-get

command | | httpd

/etc/passwd
Alias /doc/ /usr/share/doc/
<Directory /usr/share/doc>
    Options Indexes FollowSymLinks
</Directory>

sudo /sbin/chkconfig httpd on sudo /sbin/service httpd start guides

bash
shrike

§ cd /home/nfs/redhat-9/doc/RedHat/RPMS sudo rpm - Uhv *.rpm

rhl-gsg-en-9
index.html

Maximum RPM

maximum-rpm-1.0

moderation

bash
man
info
bash
pinfo
bash
bash
rpm
/usr/lib/libc.so

man bash info bash pinfo bashbash : pinfo : commandnotfound sudo rpm - Uhv